2.26 gm of a sample of dolomite containing only caco3 and mgco3 were dissolved in 700 ml of 0.1 N. The solution was diluted to 2.5 l. 25ml of this solution required 20 ml of 0.01 naoh solution for complete neutralisation. Find  %composition of the ore. 

Grade:11

1 Answers

Arun
3 years ago
equivalent of caco3+equivalent of mgco3 = 36
 
moles+moles=36/2000
 
moles+moles=0.018
 
w/100+1.64-w/84 = 0.018
 
164-16w = 151.2
 
16w=12.8
 
w=0.8 gm
 
weight of caco3 = 0.8 gm
 
weight of mgco3 = 0.84 gm
 
percentage composition of caco3 = 0.8/1.64*100 = 48.79
 
percentage composition of mgco3 = 51.21